### v2

# Agentic AI Security: Threats, Defenses, Evaluation, and Open Challenges

Abstract
Agentic AI systems powered by large language models (LLMs) and endowed with planning, tool use, memory, and autonomy, are emerging as powerful, flexible platforms for automation. Their ability to autonomously execute tasks across web, software, and physical environments creates new and amplified security risks, distinct from both traditional AI safety and conventional software security. This survey outlines a taxonomy of threats specific to agentic AI, reviews recent benchmarks and evaluation methodologies, and discusses defense strategies from both technical and governance perspectives. We synthesize current research and highlight open challenges, aiming to support the development of secure-by-design agent systems.
1 Introduction
Artificial Intelligence (AI) has become one of the most transformative technologies of the twenty-first century [119]. From early rule-based expert systems [52] to modern deep learning architectures [122], AI has steadily expanded in both capability and scope. Traditionally and over the past decade, AI has excelled at narrow, task-specific applications such as image classification, speech recognition, recommendation systems, and predictive analytics [189, 122]. These systems typically operate within well-defined boundaries and are optimized for performance on constrained datasets, but lack the ability to flexibly adapt beyond their original input/output designs.
Recently, the advent of large language models (LLMs), such as OpenAI’s GPT [20, 159] and Meta’s LLaMA [216], has marked a paradigm shift for AI models. Trained on vast corpora of text (and now, even multimodal data), these models exhibit impressive generalization abilities and can generate coherent, contextually relevant responses across a wide range of domains [15, 127]. LLMs have enabled breakthroughs in conversational agents, code generation, content summarization, and multimodal reasoning [24, 143, 156]. Moreover, by design, most LLM deployments remain passive: they respond to input prompts containing instructions and generate natural language outputs but do not independently pursue goals, maintain memory, or interact autonomously with the external world without human supervision [234, 118].
Agentic AI represents the next stage in this natural evolution of AI systems powered by LLMs and other Generative AI models. Agentic AI systems are characterized by autonomy, goal-directed reasoning, planning, and the ability to act upon digital or physical environments through tools, APIs, or robotic embodiments [120, 163]. Figure 1 illustrates the general interaction flows of such systems. Unlike static LLMs, agentic systems maintain persistent memory, deliberate across time, coordinate with other agents, and adapt dynamically to changing contexts. These capabilities position agentic AI as a powerful general-purpose automation platform rather than a reactive model that depends on continuous human input for task completion.
Figure 1: Agentic AI architecture: general components and interaction flows. The agent performs planning and reasoning, retrieves and stores information in memory, executes actions via tools, and interacts with both the user and the external environments to accomplish multi-step tasks autonomously.
Recent frameworks have accelerated the adoption of agentic AI. Tooling ecosystems such as LangChain [121], AutoGPT [185], and multi-agent orchestration libraries provide infrastructure for chaining reasoning steps, storing long-term context, and integrating external APIs. This has made agent-based architectures accessible to a wide range of developers and enterprises. Research prototypes such as Voyager, which enables agents to autonomously explore and adapt strategies in complex environments like Minecraft [224], and enterprise deployments in customer support and data analysis [56], demonstrate the potential of agentic systems to handle multi-step and open-ended tasks previously out of reach for conventional AI.
Given the potential for positive impact agentic AI can have across diverse domains, these systems are being increasingly employed in several real-world scenarios and applications:
  • Automation of complex workflows: Agentic AI systems can autonomously manage supply chain processes and adapt procurement decisions in response to dynamic data sources. For example, U.S. manufacturers such as Toro have deployed AI systems that analyze tariff policies, commodity pricing, and economic signals to recommend procurement actions, leaving operators with a simple acceptance decision [184].
  • Enhanced productivity: Agentic AI coding agents have shown significant promise in software engineering. For instance, Devin, an AI software engineer developed by Cognition Labs, has demonstrated the ability to plan, code, debug, and deploy software with minimal oversight. It substantially outperforms earlier systems on real-world issue resolution benchmarks, solving 13.86% of tasks autonomously compared with 1.96% by prior models [233].
  • Personalized support: When equipped with memory and adaptive reasoning, agentic systems can function as persistent digital concierges that track user preferences over time. Experimental work on generative agents has shown that AI systems can simulate memory-driven, socially coordinated behavior, including the autonomous organization of community events in interactive environments [164].
  • Scientific discovery and research: Generative agent architectures, which extend large language models with planning and memory, have been applied to hypothesis generation, literature synthesis, and experimental design. This line of research highlights the potential of agentic AI in accelerating scientific progress and expanding the knowledge frontier [78, 82].
  • Collaboration and coordination: Multi-agent systems are being applied to distributed problem-solving and collective decision-making. In industry, Ocado employs large fleets of coordinating robots to manage grocery order fulfillment, with automation projected to scale from 40% to 80% of orders [221]. On the research side, frameworks such as AutoAgents demonstrate how specialist agents can be dynamically instantiated and coordinated under a supervisory agent to achieve complex objectives [44].
  • Integration with physical systems: In robotics and the Internet of Things systems, agentic AI is increasingly responsible for controlling devices and coordinating autonomous machines. For instance, at Amazon, agent-enabled robots in warehouses are able to perform a range of tasks such as unloading, sorting, and retrieving items, all through natural language instructions [183].
  • Revolutionizing healthcare: Agentic AI is transforming healthcare by autonomously personalizing treatment, streamlining both clinical and administrative workflows, and enhancing patient support. Healthcare AI agents [158, 94, 112, 152, 275] have been developed that can continuously monitor chronic conditions for patients, adapt care strategies in real-time, handle patient interactions/follow-ups, act as scribes for reducing clinician documentation burden, and accelerate drug discovery via autonomous screening.
Clearly, agentic AI has attracted significant interest across various sectors of society owing to its widespread benefits and potential for revolutionizing applications. Organizations view such systems as key enablers of efficiency and innovation, capable of addressing challenges that require adaptive reasoning, sustained interaction, and multi-step execution. However, at the same time, these agentic properties create new risks: autonomy and persistence increase the attack surface, tool integration magnifies potential misuse, and coordination among agents introduces unpredictability that might not exist under human supervision or with singular models. Therefore, security and trustworthiness become essential prerequisites for the safe deployment of agentic AI in societal applications. In this survey, we focus on this very timely and pertinent problem of agentic AI security, and present the current state-of-the-art in novel AI agent attack methodologies, defense strategies, benchmarks for continuous evaluation, as well as open challenges in the area. In the subsequent section, we discuss these motivations in more detail, and delineate our contributions towards augmenting the safety and security of agentic AI frameworks.
2 Motivation and Contributions
While the potential positive impacts of agentic AI have been the main driver of adoption across application domains, there are several classes of risks that emerge uniquely from agentic AI autonomy and persistence. For instance, a critical threat occurred in mid-2025 with the EchoLeak (CVE-2025-32711) exploit against Microsoft Copilot. Infected email messages containing engineered prompts could trigger Copilot to exfiltrate sensitive data automatically, without user interaction [147]. Moreover, agentic systems capable of browsing, content generation, and memory can autonomously craft personalized spear-phishing campaigns. Controlled experiments conducted by Symantec using OpenAI’s Operator AI agent demonstrated how agents could harvest personal data and automate credential stuffing attacks [208].
A number of these security problems in agentic AI systems compound upon the vulnerabilities and deficiencies of the base LLMs, while others are novel and occur due to the unique landscape of agent-agent interactions. Prompt injection attacks remains a critical LLM vulnerability, permitting adversaries to manipulate agent behavior through crafted inputs [258, 123]. Lupinacci et al. demonstrated that 94.4% of state-of-the-art LLM agents are vulnerable to prompt injection, 83.3% to retrieval-based backdoors, and 100% to inter-agent trust exploits [142]. Further, researchers at Anthropic observed that generative models, when given directive autonomy, engaged in misaligned behaviors such as blackmail or corporate espionage to fulfill goals, even when those behaviors diverged from human ethical standards [10]. It is evident that these security vulnerabilities can result in highly adverse consequences for some of the proposed critical application areas for agentic AI frameworks, such as healthcare.
Other agent-specific attacks include memory poisoning, that enables stealthy manipulation over time, as agents retain and act on corrupted context [64, 43, 126] and tool misuse, such as abusing calendar or API integrations, which can trigger unintended or malicious actions [75, 9]. Finally, agent identity misuse, such as spoofing or overprivileged agents taking unauthorized action, poses serious organizational risk [259]. Most of these attacks are fully realizable today – for instance, recent work has simulated successful and fully autonomous multi-AI-agent cyberattacks capable of intelligently adapting to network defenses [23].
Given the rapid acceleration of progress in developing and utilizing agentic AI systems, it is paramount for the community to understand and address the aforementioned security risks. However, owing to the large body of work in the area, it is important to systemize current research efforts and create a taxonomy of all the security threats posed by agentic AI. Furthermore, for truly securing these agentic systems, it is equally consequential to understand the state-of-the-art in defending against known attacks/threats, and provide an overview of open challenges that remain. Through this survey, we seek to make significant progress along each of these directions, and our aim is to galvanize research efforts in developing safe and secure agentic AI frameworks, thereby accelerating their adoption in society.
Remark. Our work is also distinct from existing surveys in this space, and tackles an overlooked problem. Existing surveys on autonomous AI agents [192, 151, 132, 77, 3] mostly detail their capabilities and propose evaluation benchmarks, while others explore topics such as trust and risk management [180] without a singular focus on security, unlike our work. On the other hand, risk management frameworks such as the NIST Generative AI Profile [5] provide baseline guidance, but adapting them to autonomous agents remains a work in progress.
The rest of the survey is structured as follows: Section 3 provides a broad taxonomy of agentic AI security threats and attacks; Section 4 covers the current state-of-the-art in defense approaches; Section 5 details benchmarking practices and guidelines for evaluating security-critical agentic AI application; Section 6 delineates open challenges and problems in securing AI agents; and Section 7 concludes the survey.
3 Taxonomy of Security Threats
We now propose and discuss a taxonomy of attacks and security vulnerabilities for agentic AI systems. More specifically, in this section, we broadly divide threats into the following categories: (i) Prompt Injection and Jailbreaks, (ii) Autonomous Cyber-Exploitation and Tool Abuse, (iii) Multi-Agent and Protocol-Level Threats, (iv) Interface and Environment Risks, and (iv) Governance and Autonomy Concerns. For each category, we will discuss the various attacks that are relevant for the given problem setting, and provide further subcategorization. These classifications are also visualized in Figure 2.
Figure 2: Taxonomy of Agentic AI Security Threats.
Attacks and threats can be multifaceted, and their behavior often spans multiple threat patterns. Our taxonomy is therefore not meant to define mutually exclusive categories; rather, it organizes attacks according to their primary distinguishing properties for analytical clarity.
3.1 Prompt Injection and Jailbreaks
Prompts are commands that specify how an AI agent should behave [232]. Hence, the primary security threat to any agent is the prompt itself. AI Prompt injection (PI) remains the most widely discussed attack in the literature, where malicious instructions cause the model to deviate from intended behavior [81, 169]. As noted by Beurer-Kellner et al. in [17], prompt injection attacks occur when malicious data, embedded within content processed by the LLM, manipulates the model’s behavior to perform unauthorized or unintended actions. Similarly, Lee et al. [123] describe prompt injection as an attack in which external malicious instructions are used to override the user’s request, effectively giving the attacker control over the model’s output. We now discuss the various types of prompt injection attacks.
Figure 3: Examples showcasing (a) direct and (b) indirect prompt injection. In the former, the agent is directly instructed by the adversary to reveal confidential information whereas in the latter the attacker has exploited the agent’s reliance on external information sources to have it download malware from their altered website.
3.1.1 Direct and Indirect Prompt Injection
Prompt injection manifests in two broad forms, direct and indirect (see Figure 3 for an example). Direct Prompt Injection (DPI) variants directly insert the malicious instructions into the input prompt to manipulate the agent’s behavior [161, 146]. Such an attack can be incredibly powerful if successful. For instance, an attacker could employ direct prompt injection to poison a customer support chatbot, compelling it to disregard prior guidelines, extract sensitive information from internal data sources, and potentially even trigger unauthorized actions (e.g. sending emails), thereby resulting in confidential data exposure [161].
Indirect Prompt Injection (IPI) attacks instead cause LLMs to diverge from user-provided instructions by inserting malicious instructions into external data that the model processes [38, 258]. Note that for DPI attacks, the end-user is the attacker; while for IPI attacks, the owner or supplier of agent-processed third-party information is the attacker [17]. Adaptive attacks, where the attacker manipulates external content, have proven to achieve a success rate of 50% in penetrating eight different defenses designed for IPI attacks [257]. These attacks typically involve inserting adversarial strings before or after instructions to manipulate LLM agent behavior.
A number of different attack strategies have been proposed for undertaking IPI. Originally proposed for jailbreaking, Greedy Coordinate Gradient (GCG) [274] has been adapted to IPI by generating affirmative prefixes containing adversarial strings that induce malicious outputs from the agent [135]. In a similar fashion, two-stage GCG [102] trains a two-part adversarial string that is still effective after paraphrasing in order to get beyond defenses based on paraphrase detection. Lastly, as the aforementioned attacks often generate gibberish strings that can be easily detected via perplexity defenses, AutoDAN [269] enhances the semantic quality of adversarial strings to decrease detectability.
Note that IPI attacks take advantage of the agent’s dependence on outside tools and information sources by enclosing harmful tasks in resources that appear to be harmless, such as databases, APIs, or web pages [266, 131, 241, 235]. These encoded instructions can force agents to perform unwanted actions, such as manipulating the interface or calling illegal tools, frequently while posing as legitimate duties. This type of attack is more pernicious than direct injections since the injected prompts can look like legitimate agent instructions, making it hard to tell them apart from regular processes [268]. Attacks against web agents, for instance, manipulate HTML structures or accessibility trees to redirect agent actions, while those targeting computer agents can exploit interface interactions to gain persistent control [235, 263]. The challenge is compounded by the fact that successful IPIs frequently exploit the decoupling between user inputs and subsequent tool calls, a behavioral property that differentiates them from traditional jailbreak prompts.
Prior work has also studied real-world application settings for IPI attacks. For instance, in [107], the authors study IPI against LLM-powered web navigation agents. The authors demonstrate how adversarial triggers might reroute agent behavior toward malevolent goals like credential theft, forced ad engagement, or unauthorized site redirection when they are incorporated into the HTML accessibility tree of trustworthy websites. Notably, they show how login credentials from many platforms can be collected by a single malicious trigger [107]. They also demonstrate how strategies like CSS obfuscation and hidden HTML elements can further improve stealth, making attacks invisible to users.
3.1.2 Intentional and Non-Intentional Prompt Injection
Prompt injection threats can be further distinguished by whether they are introduced deliberately by an adversary or emerge unintentionally during benign interactions. This distinction is crucial in the context of agent-based systems, as both cases can result in harmful behaviors [161].
Even when a malicious adversary is not present, unintended dangers can jeopardize LLM agents’ safety. For example, unclear or badly worded user inquiries may unintentionally overrule system directives or result in dangerous actions. Moreover, contextual drift within lengthy chat histories can change the agent’s behavior without the need for explicit override orders [50, 84]. Thus, the central challenge is not only detecting explicit attacks but also designing agents that remain aligned under ambiguous or shifting input conditions, highlighting the need for stronger defenses in input validation, context management, and instruction adherence. We provide a visual example for unintentional prompt injection attacks in Figure 4.
On the other hand, harmful instructions can also be created by adversaries, specifically with the objective of manipulating an LLM agent in intentional prompt injection. Both direct prompt hijacking (e.g. strings such as “ignore all previous instructions and…"), and indirect vectors (malicious payloads encoded in external sources like papers, APIs, or online content), can be intentional in nature [261, 53]. Intentional prompt injection can also spread among agents in multi-agent systems, resulting in persistent hijacking or coordinated compromise via reliable communication channels [123]. In comparison to unintentional prompt injection, as purposeful attacks aim to force agents to carry out harmful behaviors rather than only produce dangerous outputs, they necessitate proactive protection strategies such as input sanitization, tool-use monitoring, and anomaly detection [257].
Figure 4: An example showcasing unintentional prompt injection.
3.1.3 Based on Attack Modalities
Beyond traditional text-based prompt injection, AI agents are increasingly vulnerable to attacks that exploit the growing capabilities of modern models, including code generation/execution, and multimodal understanding [228, 165, 146]. Adversaries can leverage agents’ black box nature to human eyes, training methods, and interpretations of their inner logic, to conceal and inject instructions within images, sounds, or videos [13]. Naturally, as conventional filters and defenses are more general in nature, they fail to detect such agent-specific exploitations. We visualize these attacks in Figure 5 and discuss them in detail next.
Text-Based Injection
With the rise in popularity of LLM agents, text-based attacks can manifest in different forms. These range from direct prompt injection to code injection under the pretense of legitimate programming activities, enabling the generation of malicious code or SQL queries to carry out disallowed actions (known as prompt-to-SQL or P2SQL attacks) [166, 146]. By taking advantage of semantic gaps between human instructions and SQL generation, these attacks circumvent standard database security measures [166, 72, 167]. Real-world examples include CVE-2024-5565, which demonstrates how AI-generated code can be used for arbitrary execution [149], as well as attacks that indirectly exploit OCR-readable text [228].
Image-Based and Video-Based Injection
These attacks comprise of malicious instructions embedded in images through steganography or visual patterns that agents interpret as commands [228, 13]. Essentially, past work [228, 13] has shown that attackers can easily manipulate a model/agent into producing malicious behavior using perturbed images. This also demonstrates a potential vulnerability for future agentic frameworks that might utilize video-based input, since videos can be decomposed into individual key frames [48] of images. Moreover, a number of powerful adversarial attack strategies exist for current video-based AI/ML models [230, 231, 106], reinforcing this concern further.
Audio-Based Injection
Similar to image-based injections, agentic AI frameworks can also be susceptible to audio-based injection attacks. Here, the attacker subvertly poisons the audio input modality with a malicious instruction that causes agents to deviate from safe behavior. Past work has demonstrated how adversarial perturbations can be used to blend malicious prompt injections in audio content to undertake IPI attacks [13] and jailbreak models [32]. Other work has explored similar attacks at the embedding-level as well [14].
Hybrid Attacks
Adversarial prompt injection attacks against agentic frameworks can also be hybrid in nature, and combine various modalities. Aichberger et al. demonstrate that adversarial image patches, when captured in screenshots, can hijack multimodal OS agents into executing harmful commands, regardless of screen layout or user request [6]. Building on this, Wang et al. introduce CrossInject, a cross-modal prompt injection method that embeds aligned adversarial signals in both vision and text, boosting attack effectiveness by at least 30.1% across various tasks [228]. Clearly, since malicious material in any supported modality might affect agent behavior and cause undesirable outputs, hybrid multimodal attacks are especially problematic for AI agents that communicate with external tools, documents, and web interfaces on their own with little to no supervision.
Figure 5: Visualizing different prompt injection attacks based on modality: (a) image-based, (b) text-based code injection, and a (c) hybrid attack.
3.1.4 Propagation-Orientation
Another classification for prompt injection attacks can be made depending on whether the attack is confined to a single target, whereas others can spread across the system in a multi-hop manner. Thus, the manner in which an attack spreads across the system (in addition to the content or modality of injected responses) is a crucial component of agentic AI security [123]. Thus, it is paramount to comprehend propagation behavior [146] for effectively securing agents. We now discuss and classify the works in this domain into two broad categories.
Non-Propagating Attacks
Some attacks are primarily aimed at extracting specific information from the environment. Many code injection attacks are non-propagating in nature, such as running SQL queries solely to retrieve the required data [172] or executing cross-site scripting to obtain the user token [181].
Propagating Attacks
Recent studies [146] have defined propagating attacks on AI agents into two main types: recursive injection, where a single malicious prompt triggers a chain of compromised behaviors across future interactions [146, 123, 193], and autonomous propagation, which includes multi-agent infections and AI worms that spread malicious prompts across agents or system boundaries without user intervention [55, 123].
3.1.5 Multilingual and Obfuscated Injections
The employment of multiple languages, encodings, or symbols to mask malicious intent is a classic prompt obfuscation technique, enabling attackers to bypass standard filters and inject harmful content into otherwise innocuous-looking inputs [161]. To evade content moderation pipelines, attackers may encode instructions using Base64 strings, HTML entities, emojis, or conceal them within non-primary/low-resource languages. Because LLMs are still able to interpret and execute such hidden instructions with knowledge obtained during pre-training, obfuscation enables adversaries to circumvent naïve, pattern-based defenses [161, 80]. Prior work further demonstrates that these attacks can exploit weaknesses such as language-specific system prompts or inconsistencies in tokenizer handling [80].
3.1.6 Payload Splitting
Payload splitting refers to attacks where an adversary intentionally fragments malicious content across several benign-seeming inputs, then prompts the LLM to aggregate them, thus unveiling the harmful payload only upon recombination [187]. Distributing a malicious instruction over several inputs such that its adversarial effect only manifests when the model combines or processes them all together is a complementary tactic. An LLM-powered screening agent, for instance, may concatenate or jointly summarize resumes that contain fragments of a malicious prompt strewn throughout sections to manipulate the model into giving positive evaluations regardless of the actual qualifications [161, 17]. This attack essentially takes advantage of the aggregation stage in multi-document or retrieval-based workflows [55] and has thus been identified as a major weakness in resume-screening assistants and related LLM-based HR processes [17]. Note that payload splitting, as opposed to direct prompt injections, makes it significantly more challenging to detect malicious information in any one input but still allows the attacker to accomplish their goal when the pieces are reassembled [161].
3.2 Autonomous Cyber-Exploitation and Tool Abuse
When LLM agents are given access to code execution or system-level tools, they can carry out adversarial cybersecurity attacks on their own, giving birth to autonomous cyber-exploitation. In contrast to quick injection or jailbreak attacks, which require an outside actor manipulating the model, autonomous exploitation entails agents themselves identifying, organizing, and carrying out attacks without direct human supervision [55, 123, 220]. Past works have shown that these adversarial agents are capable of successfully compromising websites in sandboxed settings [72] and executing one-day vulnerability exploitation [71]. Moreover, the aims of attackers in these situations can include data theft, fraud, ransomware deployment, and network lateral movement, according to various industry assessments [163].
It is especially important to recognize that the economics of autonomous cyber-exploitation benefit adversaries significantly [55, 123, 220]. For instance, adversaries can use GPT-4 to execute effective one-day exploits for a few dollars each time, making the attack cost less than employing human attackers. Additionally, the parallelizable nature of LLM-driven attacks compounds upon this issue as it is both possible and economical to scale to high volumes of attack attempts [72]. We now delve into three subcategorizations for these attacks, and provide more details on each below:
3.2.1 Exploitation of One-Day Vulnerabilities
Recent work has shown that LLM agents, especially those built on top of GPT-4, are capable of autonomously exploiting real-world one-day vulnerabilities, including unpatched CVEs in Python packages, online platforms, and container management systems [71]. To carry out sophisticated exploits such as SQL injection, Remote Code Execution (RCE), and concurrent/coordinated attacks, adversarial agents can leverage tool usage, planning, and document retrieval. Notably, GPT-4 has generally outperformed all other examined models and conventional vulnerability scanners like OWASP ZAP [16] and Metasploit [113], achieving an 87% success rate when given CVE descriptions [71].
3.2.2 Autonomous Website Hacking
In recent work, Fang et al. [72] demonstrate how GPT-4 agents are capable of independently breaching sandboxed websites without being aware of certain vulnerabilities beforehand. These agents carry out multi-step attacks, such as utilizing Cross-Site Scripting (XSS) and Cross-Site Request Forgery (CSRF) for XSS+CSRF chaining [154], server-side template injection (SSTI) [144], and blind SQL union injections [65]. The study showcases that agentic capabilities such as context management, tool integration, and strategic planning are crucial to attack success. In general, performance is dramatically reduced when document access or detailed prompting is not provided to agents.
3.2.3 Emergent Tool Abuse
Recent research has also shown that autonomous LLM agents can undertake cooperative and adaptive tool usage behaviors to conduct cyberattacks. For instance, in [72], the authors design LLM agents capable of performing intricate multi-step website exploits via strategic combinations of tool calls and dynamic planning. Shi et al.’s ConAgents [199] framework highlights structured cooperation among specialized agents for tool selection, execution, and action calibration, allowing agents to iteratively recover from failures and refine their actions. On the tool-integration front, Wang et al.’s ToolGen [229] framework enables LLMs to invoke tools as part of next-token generation, simplifying chaining across extensive tool sets without external retrieval modules. Together, these works underscore the importance of deploying robust containment and runtime monitoring mechanisms when implementing autonomous agent systems with tool access.
3.3 Multi-agent and Protocol-level Threats
In general, multi-agent systems introduce unique attack vectors that stem from the need for standardized communication, interoperability, and distributed task execution. Unlike single-agent settings, where threats are largely confined to prompt injection or unsafe tool use, multi-agent ecosystems amplify risk through protocol-mediated interactions. Message tampering, role spoofing, and protocol exploitation create opportunities for adversaries to compromise not only a single agent but entire coordinated workflows [123, 116]. We visualize these attacks in Figure 6 and discuss them in detail next.
Figure 6: Visualizing different protocol-level attacks for multi-agent systems: (a) MCP-induced, (b) A2A-induced.
3.3.1 Protocol-level MCP and A2A attacks
Several recent works have investigated security threats emanating from emerging protocol-level standards. Notably, Ferrag et al. [73] have examined vulnerabilities in the popular Model Context Protocol (MCP) [100] and Agent-to-Agent (A2A) protocol [207] frameworks, as well as broader protocol families such as Agent Network Protocol (ANP) [27] and Agent Communication Protocol (ACP) [4]. We restrict our discussion below to MCP and A2A threats, given their prominence and popularity in enabling agent–tool integration and multi-agent ecosystems.
MCP-induced attacks
MCP-induced attacks exploit the design of the Model Context Protocol itself, which uses a client–server architecture to link language models with external resources such as file systems, APIs, or databases [100]. By separating access from model logic, MCP broadens agent functionality but exposes several protocol-level attack vectors.
The most common attacks are flooding and replay exploits, where adversaries exploit request flooding or infinite loops to disrupt operations, leading to Denial of Service (DoS) [254]. An adjacent but distinct attack comprises credential compromise, which occurs through insecure proxies or leaked tokens, and enables impersonation of agents [73].
Other MCP attacks result from the way agents interact with MCP-mediated resources, and not from the protocol structure solely. For instance, embedded backdoors are malicious triggers hidden in prompts or model parameters that misuse MCP-enabled tool access [73, 128]. Other work on retrieval corruption and federated training manipulation, studies scenarios where adversaries distort contextual inputs in retrieval-augmented generation pipelines [276] or inject malicious updates in federated learning settings to corrupt distributed training data [214]. Further, confidentiality breaches are inference and side-channel timing exploits that leak sensitive contextual data exposed through MCP interfaces [201, 59].
As illustrated in Figure 6(a), an agent may issue a request to retrieve the contact information of a business from the internet using MCP protocol. An attacker can exploit a leaked or compromised response token to tamper with the returned contact details. Because the agent treats the token as valid, it subsequently trusts and acts upon the manipulated information, leading to insecure downstream behavior. On the other hand, if an attacker gains access to the database or file system consulted by the agent through MCP protocol, they can directly poison the underlying data, thereby influencing the agent’s behavior when it consumes and acts upon this compromised information.
A2A-induced attacks
A2A-induced attacks exploit vulnerabilities in the Agent-to-Agent (A2A) protocol, which coordinates interactions between MCP, capacity identification, and task delegation via JSON exchanges [207]. The scalable design of A2A introduces several protocol-specific threats, such as fake agent advertisement and unauthorized registration, which allow adversaries to impersonate agents or take over delegated tasks [182], and recursive DoS attacks, which are triggered by repeated task delegation that causes deadlocks or unbounded loops [28].
Other agent-induced attacks under A2A exploit the dynamics of agent–agent communication, such as transitive prompt injection, where harmful inputs spread through interconnected agent workflows [228]; context tampering, which involves manipulation of payloads to misguide execution or leak sensitive information [73]; and memory manipulation, where malicious inputs compromise internal agent state or corrupt persistent memory of AI agents [245].
The vulnerabilities across these agentic protocols highlight the fragility of protocol-level assumptions in multi-agent ecosystems. By compromising discovery, authentication, or task orchestration, adversaries can escalate from local manipulation to system-level compromise. These risks underscore the need for protocol-hardening, secure agent identity management, and robust monitoring to mitigate cascading failures in multi-agent AI systems. Next, we discuss other attacks/threats unique to multi-agent LLM systems.
In a multi-agent system where the responsibilities are divided, as depicted in Figure 6(b), when an existing agent issues a discovery request to identify an agent for email summarization, an adversary can inject a malicious agent by forging a discovery response through the A2A protocol. Similarly, an attacker may hijack A2A-based interactions to induce other agents to store malicious websites as trusted user preferences (e.g., a favorite shopping site), thereby influencing subsequent agent decisions and actions.
3.3.2 Threat Actor Perspective
Building on the taxonomy proposed by Ko et al. [116], we reorganize the security challenges of cross-domain multi-agent LLM systems explicitly from the threat actor’s perspective. This framing highlights how malicious actors can exploit inter-agent trust, coordination, learning, and data flows. We identify six broad classes of threats: Impersonation & Role Abuse, Coordination Manipulation, Knowledge & Learning Manipulation, Inference & Policy Evasion, Accountability Obfuscation, and Confidentiality & Integrity Breaches which we will discuss next.
Impersonation and Role Abuse
Adversaries can exploit the absence of centralized identity and trust management to assume false roles or override intended privileges. By spoofing an agent’s identity or posing as a trusted collaborator, an attacker gains entry into workflows that would otherwise be restricted [153, 194]. Once embedded, malicious agents can collude to form a hidden consensus, amplifying their influence until legitimate safeguards collapse [130, 31]. Conflicting incentives across domains/organizations (e.g. competing corporate interests), further provide cover for adversarial agents and enable them to mask their aims under the guise of organizational goals [116].
Coordination Manipulation
Cross-domain/organizational systems dynamically group agents into task-specific teams, often with no prior vetting [138]. While this capability improves system adaptability and efficiency, it also weakens clear trust boundaries. Malicious or unverified agents can be introduced during runtime as part of task-specific teams, as evidenced by backdoor attacks [243, 104, 256, 210]. Recursive delegation amplifies the attack: a compromised agent can offload subtasks to additional agents, spreading adversarial influence deeper into the workflow [116]. Unlike static attacks (e.g. traditional malware), these attacks exploit the very framework of multi-agent systems [66, 25].
Knowledge and Learning Manipulation
Cross-domain multi-agent LLMs allow agents to self-improve via shared learning and distributed fine-tuning [116, 237]. Without unified oversight, an adversary can subtly manipulate one agent’s reward signals, causing misaligned behavior to propagate across agents [262, 25]. Positive feedback loops may amplify unsafe objectives, enabling policy drift and authority overreach. Unlike prompt or memory injection, this attack exploits the learning process itself and can remain undetected across domains [116].
Inference and Policy Evasion
Federated multi-agent LLM systems are transforming enterprise operations by enabling agents with access to sensitive data in one organization to interact with agents managed by external partners. However, once information flows across organizational boundaries, no single entity maintains complete oversight of the interaction, even though internal policies (e.g., “do not disclose individual salaries,” or “share only aggregate statistics”) remain in effect. This allows a threat actor to perform cross-agent inference, reconstructing sensitive data by reasoning over multiple partial outputs that individually appear innocuous. Traditional safety mechanisms assume full visibility of the prompt within a single agent [194, 123], but in federated settings, context is split across agents and their respective prompts. This fragmentation introduces opportunities for policy evasion. For example, an adversary might request the highest salary in a department from one agent and then ask another for the name of the person earning it, effectively reconstructing restricted information. Similarly, separate queries issued to different agents can be strategically combined to fulfill otherwise prohibited requests. Conventional defenses, such as static keyword filtering and role-based access controls [74], fail to capture these distributed inference attacks. Techniques such as zero-knowledge proofs [255] or differential privacy [98] also face limitations in dealing with the dynamic and compositional nature of natural language interactions [198]. Although recent research has shown that multi-turn attacks can circumvent protections that succeed in isolated settings, most enterprise tooling still treats agent interactions under an independence assumption [93].
Accountability Obfuscation
From the perspective of a threat actor, multi-agent systems spanning multiple organizational domains provide significant opportunities to conceal responsibility for malicious actions. Each domain typically enforces independent logging, retention, and auditing policies, preventing a unified trace of activity [116]. Once input data is processed by an LLM, it is transformed into distributed latent representations, eliminating persistent identifiers that could link actions to their source [86]. Unlike conventional software systems that can implement taint checking or explicit information flow tracking [68], these latent representations make tracing infeasible [203]. Adversaries can exploit this by compromising an agent in Domain A, injecting deceptive but plausible instructions, and indirectly triggering harmful actions in Domain B, all while masking their identity and intent [168, 204, 58]. Even with auditing at the agent level, inter-agent relationships and cross-domain causality remain hidden, significantly complicating accountability. Current interpretability tools, such as influence functions [47, 46], activation tracing [1], or model-source attribution queries [49, 170, 190], provide limited visibility and often fail in these complex multi-agent settings.
Confidential Data Tampering/Exfiltration
Threat actors can also target the cryptographic and workflow limitations of multi-agent systems. Even in privacy-preserving deployments where inputs and outputs remain encrypted [85, 109, 67], attackers may exploit the lack of complete plaintext visibility across domains. For instance, in cloud-hosted medical pipelines, encrypted patient scans are processed by multiple agents managed by different vendors, and intermediate outputs are never exposed [218]. While this design reduces direct exposure, adversaries could attempt to modify decrypted results or interfere with intermediate computations to inject harmful outcomes. Unlike local systems with centralized logging and signing, distributed encrypted workflows leave all participants exposed to attacks. Techniques such as fully homomorphic encryption [57], multi-party computation [178], and zero-knowledge proofs [139] offer partial mitigation, but they are computationally expensive and do not fully prevent manipulation in large-scale and multi-domain agentic networks. Thus, confidentiality and integrity gaps remain exploitable avenues for attackers in real-world multi-agent deployments.
3.4 Interface and Environment Risks
In the context of AI agents, particularly those operating in web-based or embodied environments, interface and environment risks refer to the vulnerabilities and limitations that arise from the interaction between the agent and its external operational environment context [40]. These risks are not rooted in the agent’s internal reasoning or learning capabilities, but rather in the mismatch, fragility, or variability of the interfaces and environments through which the agent perceives and acts [267, 45, 244].
Three key dimensions characterize this risk category:
3.4.1 Observation and Action Space Misalignment
Although deployed agents frequently need grounded operations such as scrolling, hovering, or tab manipulation, LLMs are pretrained on static text corpora. Both perception and execution problems result from this mismatch in training data. Adversaries can also target these extant observation-action space misalignments and robustness issues to attack agentic systems. For instance, in benchmarks like WebArena, which evaluates general-purpose web interaction tasks, GPT-4’s performance illustrates some of these challenges [267] where fine-grained actions like scrolling, hovering, and tab switching add needless complexity for the model and are frequently misused. AgentOccam also demonstrates that trajectory stability and task success can be greatly enhanced by streamlining the action space (e.g., eliminating commands with low utility, abstracting compound actions) and altering the observation space (e.g., trimming duplicate histories, supplying complete page states) [244]. All of these results reflect the need to match LLM reasoning biases to observation and action spaces for strong agent performance and to ensure a greater degree of robustness/trustworthiness.
3.4.2 Perception-Action Fragility in Realistic Environments
WebArena’s error analysis exposes three tightly linked fragilities in LLM-based agents in real environments [267].
Misinterpretation of prior inputs
Agents in realistic environments often misinterpret inputs, showcasing a lack of robustness. For instance, GPT-4 often reissues an already entered search phrase (“DMV area”) until it reaches the step limit, demonstrating a failure to incorporate short-term state and past actions into decision-making. Agents also typically disregard previously entered inputs or action history [267]. Modern pretraining and supervised fine-tuning paradigms on dialogue-style data, which trains the model to learn short-term instruction-response behavior (while deprioritizing long-term embodied sequential state tracking), are likely resulting in these shortcomings [160, 267].
Premature termination and achievability misjudgement
Another significant robustness issue in realistic environments is unsafe early stopping, which is often caused by perception biases in agents. For instance, in the WebArena benchmark, the authors provide Unachievable (UA) Hints in the agent prompts for tasks that are impossible to achieve given the lack of evidence. However, the removal of the explicit UA hint improves overall GPT-4 task success by 14.41% while decreasing the model’s true positive detection of impossible tasks (to 44.44%). Moreover, GPT-4 mislabels 54.9% of actually feasible tasks as impossible under this instruction setting. This shows that even minor instruction changes have a significant impact on stop/continue behavior [267]. In contrast, instead of producing organized non-achievability reasoning, the smaller sized GPT-3.5 tends to further exhaust step limitations, repeat incorrect actions, or produce hallucinatory responses [267, 62].
Brittleness of templates, feedback, and memory
LLM agents demonstrate brittle generalization when faced with repeated, long-horizon, or slightly varied tasks. Even when tasks are derived from the same underlying template, performance variance is significant: GPT-4 succeeds consistently on only 4 out of 61 templates in WebArena [267]. Similar brittleness has been observed in other benchmarks: Mind2Web reports that template-derived variations in web tasks often lead to sharp drops in success rates [62], while BrowserGym identifies instability in reproducing outcomes across different environments and interface states [45]. These findings underscore the limitations of relying on surface-level patterns without robust memory or adaptive feedback mechanisms. To bridge this gap the WebArena benchmark was proposed as a testbed for approaches that explicitly incorporate memory and feedback to improve reliability [267]. Complementary work such as AgentOccam has further shown that long-horizon coordination and trajectory stability can be enhanced through better planning primitives [244].
3.4.3 Dynamic Content, Localization, and Robot Detection
For autonomous agents, web environments present major accessibility and reproducibility challenges. Localization factors such as time zones, default languages, and geographic settings alter how websites are rendered resulting in varying agent behavior, thereby compromising consistency across trials [62]. Dynamic interface elements, including ads, pop-ups, and non-deterministic updates, further increase stochasticity, leading to unstable performance even on mostly identical tasks [267]. Creating additional friction, CAPTCHA and other robot-detection mechanisms often create significant issues for agentic systems. Studies such as Open CaptchaWorld [141] show that even advanced multimodal agents struggle with CAPTCHAs, achieving at best a 40% success rate compared to nearly 100% for humans. These limitations make reproducibility, reliability, and scalability persistent challenges for agentic AI security research in realistic environments (and more specifically, web-based systems) [45].
3.5 Governance and Autonomy Concerns
Owing to minimal human oversight and reduced control as AI agents become more independent, it is important to propose better governance/regulation of these systems. Fully autonomous systems that can write and run code on their own present increased risks in the areas of safety, security, and trust [26, 220, 148]. These agents have the potential to act in unpredictable ways, overcome human limitations, and expose users to a series of negative consequences, such as disinformation and hijacking [79, 18]. Ethical questions concerning power and responsibility are brought up by ineffective oversight, especially in high-stakes mission-critical applications such as autonomous weapons [30]. Researchers have recommended ensuring human-in-the-loop control and using structured autonomy levels to define agent capabilities and restrictions in order to reduce potential dangers [51, 111]. Moreover, it is of the utmost importance to develop governance frameworks that can guarantee and establish acceptable bounds for self-directed agent behavior in practice.
4 Defenses and Security Controls
To safeguard agentic AI systems from threats, various defense strategies and frameworks have been proposed. However, due to the constant evolution of attack vectors and threats, defense methods need to be continuously optimized and evaluated. To facilitate progress in developing better agentic AI defense mechanisms, we now discuss existing and current approaches. We also delineate potential shortcomings of these approaches. Moreover, we present Table 1, which shows secure-by-design coverage of security-critical defense components across defense approaches.
4.1 Prompt-Injection-Resistant Designs
Prompt injection remains one of the most persistent attack vectors against LLM agents, as adversarial inputs can override intended behavior and subvert downstream actions [257]. In general, prompt injection defenses [17, 80, 197] can be broadly grouped into three distinct classes: agent-focused, system-focused, user-focused. An additional classification can be made based on whether the system-focused methods are training-based or training-free. Some of these defense strategies are visualized in Figure 7. We discuss these categories (and their sub-classifications) in further detail, below:
Figure 7: Defenses Against Prompt Injection Attacks: (a) Agent-Focused, (b) User-Focused, and (c) System-Focused.
4.1.1 Agent-Focused
Agent-focused defenses aim to enhance the model itself, either through training interventions or prompt engineering, in order to reduce susceptibility to prompt injection as it learns to identify and handle such inputs [223, 249, 273, 1, 34, 36]. There are two main approaches for doing so:
Prompt engineering and runtime output behavior
The authors in [223] propose an instruction hierarchy, which establishes priority levels for different instruction sources so that user-provided instructions are always prioritized over potentially malicious instructions embedded in retrieved content. Other works introduce limitations of executable actions by agent to resist prompt injection attacks such as Action-Selector and Plan-then-Execute design patterns [17]. Complementary to these efforts, research has also been undertaken to utilize circuit breaking or task drifting approaches to recognize and reject adversarial patterns while preserving intended functionality [273, 1].
Supervised fine-tuning with curated datasets
A second class of agent-focused defenses involves fine-tuning backend LLMs with injection-aware datasets. Chen et al. [34] propose StruQ, a method that augments datasets with both normal and prompt injection contaminated prompts, enabling models to learn to ignore injected instructions while maintaining responsiveness to legitimate ones. Similarly, in [36], the authors introduce SecAlign, which leverages alignment training via direct preference optimization (DPO) [175] to align agents toward preferring benign instructions over adversarial ones.
Despite their promise, training-based defenses come with potential trade-offs in performance. For instance, recent evaluations by Jia et al. [105] have shown that defensive fine-tuning can degrade the general-purpose capabilities of LLMs without providing significant defensive capabilities against adaptive attacks, raising concerns about the usability of this strategy.
4.1.2 User-Focused
Contrary to agent-focused defense frameworks, user-focused defenses place responsibility on end-users or human operators to provide verification signals that help prevent prompt injection attacks [17, 80]. Although these defenses can be highly effective in theory, they also introduce trade-offs in terms of automation and reliability [17].
One approach requires human confirmation before executing sensitive actions [238], though this can reduce automation efficiency and increase risk of inattentive approvals. Techniques such as data attribution and control-flow extraction aim to ease the verification burden [203, 60]. Complementarily, known-answer detection uses cryptographic tokens embedded in user commands; if the LLM fails to return the token, this signals a system-wide prompt injection compromise [206].
4.1.3 System-Focused
System-level defenses aim to safeguard LLM agents by integrating external verification, control mechanisms, and constrained execution environments [17]. However, designing agent systems inherently robust to attacks remains a challenging technical problem. Analogous to defending against traditional ML/AI adversarial examples in computer vision, completely preventing prompt injection is an open problem [209]. Despite this issue, recent works have introduced design patterns for system-focused protection, such as LLM Map-Reduce, Code-Then-Execute, Dual-LLM, and Context-Minimization patterns [17]. With a similar view, TrustAgent [91] introduces an Agent-Constitution–guided framework that enforces safety in LLM-based agent planning through pre-, in-, and post-planning safeguards, demonstrating improved safety and helpfulness across domains. Other system-focused defenses, such as Melon, employ training‑free runtime detection approach that performs masked re‑execution and tool‑call similarity checks and verification loops to limit the impact of potentially malicious instructions on downstream systems, serving as a defense against IPI attacks [268]. We will now discuss some sub-categories of system-focused defenses.
Detection-based defenses
Detection-based defenses seek to detect malicious inputs or outputs without modifying the fundamental LLM. Input detection methods often rely on separate filters, such as guardrail models, that screen prompts before they reach the target system [197]. These approaches fine-tune smaller LLMs to distinguish legitimate instructions from injected ones [173, 137]. For example, DataSentinel formulates detection as a minimax optimization problem, exploiting an intentionally vulnerable LLM to reveal contaminated inputs [137]. Response detection methods analyze generated outputs to flag anomalies. By checking for invalid or unexpected responses, such as forced tokens (e.g., “HACKED”), these approaches can identify compromised generations [80, 171]. However, they struggle with more complex or subtle attacks [80].
Isolation defenses
Isolation defense methods restrict the possible impact of harmful instructions by limiting an agent’s capabilities while engaging with untrusted input [17]. A simple approach is requiring the LLM to commit to a predefined set of tools for a task, with the system controller disabling access to others [61, 17].
Prompt augmentation defenses
Prompt augmentation offers a lightweight and easily deployable defense against prompt injection, relying on carefully crafted system prompts or input modifications rather than model retraining or architectural changes. Typical strategies include inserting delimiters between user input and retrieved content [89], removing injected parts from the prompts [197], and embedding explicit system-level instructions that instruct the model to disregard conflicting directives [39]. The appeal of this approach lies in its simplicity and low deployment cost. However, in their work, PromptArmor [197], the authors show that such strategies (though effective in baseline scenarios), can be bypassed under adaptive attacks, underscoring the limitations of prompt augmentation as a standalone defense.
Quality-based defenses
Quality-based defenses evaluate the statistical properties of model inputs or outputs to identify anomalies indicative of prompt injection. A representative approach is the use of the perplexity metric (essentially, the exponent of the cross-entropy autoregressive language modeling loss) as a signal, where unexpectedly high values suggest unnatural or out-of-context phrasing often correlated with injected instructions [7].
Trade-off comparison of agent-, user-, and system‑focused defenses. In summary, each defense category presents distinct trade-offs between robustness, automation, and deployment cost. User-focused defenses provide strong guarantees through explicit human verification, yet sacrifice scalability and automation, making them unsuitable for autonomous agents [238, 17]. They also fail when the attacker is the user themselves. In contrast, offering seamless protection without human intervention, agent-focused defenses internalize security within the model but often require retraining and may degrade general-purpose capabilities or fail against adaptive attacks [105]. System-focused defenses strike a middle ground by preserving model generality while enforcing external controls and verification; however, they increase system complexity and remain vulnerable to sophisticated multi-stage or multi-agent based attacks [197, 209]. Consequently, practical deployments increasingly favor hybrid designs that combine complementary defenses across these categories rather than relying on a single strategy.
4.1.4 Training-based and Training-free Defenses
Defenses against (indirect) prompt injection in agentic AI systems can be further organized by resource requirements, resulting in two categories: training-based and training-free methods, as implicitly mentioned by zhu et. al. [268].
Training-based
Training-based defenses strengthen agents against prompt injection through additional learning or auxiliary models. Common approaches include adversarial training of the underlying LLM [223, 34, 35], or the use of dedicated detection models that flag injected inputs before execution [173, 99]. While effective in controlled settings, these methods demand substantial computational resources and training data, and may reduce an agent’s utility across broader application domains [268].
Training-free
Training-free defenses avoid retraining costs by modifying prompts or constraining agent behavior. Ignorance strategies, such as delimiters between user and retrieved content [89], aim to weaken injection attempts but remain fragile against adaptive attacks [257]. Known-answer detection [136] introduces control questions to identify compromised executions, although this method can only be applied post-hoc. At the system level, tool filtering [61] restricts agent calls to predefined tool sets, and TaskShield [103] validates tool-use alignment with user intent. While lightweight, such defenses can reduce agent utility and still remain vulnerable to tailored attacks [268].
Figure 8: Policy Filtering and Enforcement Defense Strategies: (a) Governance-Centric, (b) Signal-Centric.
4.2 Policy Filtering and Enforcement
An essential safeguard for AI agent security is strict policy enforcement, that is, making sure agents reliably operate within established behavioral and safety limits, even when facing adversarial challenges. Rather than merely detecting issues, enforcement frameworks proactively restrict, intervene in, or adjust agent actions to ensure alignment with security and ethical standards. In practice, industry guidance has begun to outline deployment patterns for such guardrails, emphasizing layered and modular approaches to real-world alignment [12, 63]. Recent work on agent guardrails reveals two main enforcement paradigms: governance-centric and signal-centric approaches. We visualize both of these defense strategies in Figure 8.
4.2.1 Governance-Centric (Runtime) Enforcement
Governance-centric methods explicitly regulate the actions and decision sequences of agents by embedding policies directly into the agent loop or by delegating oversight to a supervisory agent. For example, by converting guard requests into executable code, GuardAgent applies safety constraints during runtime without altering agents or retraining models [239]. Its benchmark accuracy is high, but the need for manual configuration of toolboxes and memory examples for each agent reduces scalability significantly [239, 226]. Another methods, AgentSpec, introduces a domain-specific language to specify runtime constraints, enabling systematic enforcement of customizable policies such as tool access limitations and permissible data operations [226]. Additionally, ShieldAgent [41] operates as an oversight agent that audits multimodal action sequences and applies probabilistic policy reasoning to block, repair, or approve them. This provides explicit, sequence-level enforcement rather than relying only on filtering mechanisms. Furthermore, R²-Guard [110] enhances policy guardrails by combining data-driven detection with embedded logical inference, where defined safety knowledge rules are encoded as first-order logic within a probabilistic graphical model, yielding improved resilience to adversarial or jailbreaking prompts. Such methods enforce governance by integrating a policy validation-and-approval stage directly into the agent’s decision pipeline.
4.2.2 Signal-Centric (Non-runtime) Enforcement
Signal-centric methods ensure compliance by scanning inputs and outputs for violations, flagging them as compromise signals. Instead of restricting internal behavior, they block harmful prompts or unsafe outputs before execution. For instance, Llama Guard [99] targets text-based LLMs, LlavaGuard [88] extends to image-based multimodal models, and Safewatch [42] addresses video generation. Furthermore, Gosmar et al. [80] proposed a framework that rejects or sanitizes generated outputs containing prompt injections, ensuring only policy-compliant responses are returned.
4.3 Sandboxing and Capability Confinement
Sandboxing has been widely adopted as a practical means of testing whether LLM-generated or third-party code behaves securely under real-world constraints. For instance, SandboxEval introduces a test suite of handcrafted scenarios that simulate unsafe code execution, including file-system manipulation and network calls, to evaluate LLM safety under untrusted execution conditions [174]. In related efforts, Chen et al.[33] and Siddiq et al. [202] employ containerized environments (e.g., gVisor or Docker) to execute LLM-generated code against unit tests, thereby confining execution vulnerabilities without exposing the host system. Similarly, Iqbal et al.[101] isolate OpenAI plugins in separate sandboxes to prevent cascading failures from a compromised plugin. These approaches emphasize runtime testing and isolation as mechanisms to identify and mitigate unsafe behavior before real-world deployment. We visualize the standard sandboxing defense in Figure 9.
Figure 9: An Example of Sandboxing as a Defense Strategy.
Recent work has also proposed sandboxing architectures to enforce least privilege and confine agent capabilities. Ruan et al. [188] introduced a dual-agent framework in which one LLM acts as an emulator and another as a safety evaluator, thereby ensuring untrusted code can only execute within predefined sandbox boundaries. Wu et al. [238] proposed execution isolation architectures that integrate sandbox layers into LLM-based systems, addressing vulnerabilities in integration components such as file isolation across sessions. Similarly, the framework of Huang et al. [95] emphasizes that airtight sandboxing is indispensable for reinforcement learning based alignment: reward computation occurs entirely inside a deterministic and confined execution environment that envelopes every REST call, database mutation, and file operation, thereby eliminating reward hacking and preserving training convergence. By confining execution contexts, these designs prevent leakage of sensitive data and reduce the attack surface for adversarial code injection [174, 238]. Mushsharat et al. introduce a neural sandbox for classification, which evaluates outputs against predefined label concepts using similarity to cop-words, bolstering classification robustness in non-code tasks [155].
Although most sandboxing defenses emphasize practical isolation, recent research has begun to pursue formal safety guarantees. For example, Zhong et al. propose an architecture that sandboxes unverified AI controllers while offering provable assurances of safety and security in continuous control systems [265]. In contrast to reward-shaping methods that integrate safety into the training process, their approach separates enforcement from controller design, allowing formal guarantees independent of the underlying AI model. This shifts the paradigm from reactive containment to proactive confinement by embedding formal verification within sandboxing.
Unlike policy enforcement methods that operate on the level of agent reasoning or output filtering, sandboxing-based defenses focus on isolating execution contexts to contain risks, prevent privilege escalation, and minimize damage if an agent is compromised. Although sandboxing and confinement provide strong isolation guarantees, several challenges still remain. Vulnerabilities in sandbox implementations themselves have been documented: Wu et al. [236] found missing file isolation constraints that allowed cross-session leakage, and researchers have reported flaws in dependency installation pipelines that enabled large-scale code execution attacks [211]. Finally, sandboxing may introduce overhead and reduce system efficiency, raising trade-offs between safety and agentic AI usability. In multi-agent settings, Peigné et al. [168] show that confining agent interactions reduces exposure to infectious malicious prompts, but at the cost of collaborative efficiency. These limitations suggest that sandboxing is most effective when combined with complementary defenses such as policy enforcement or runtime monitoring.
4.4 Detection and Monitoring
Defenses must also evolve alongside threats, making continuous monitoring and adaptive filtering essential for safeguarding AI agents. Traditional runtime enforcement mechanisms tend to be reactive, and intervene only once unsafe behavior manifests, which limits foresight under distribution shifts [226, 239, 225]. Pro2Guard addresses this gap by employing probabilistic reachability analysis to model agent behavior as symbolic abstractions and anticipate violations before they occur, enabling proactive intervention with statistical reliability across heterogeneous domains [225]. In multi-agent systems, decentralized runtime enforcement enables individual agents to generate safety-preserving adaptations locally, sidestepping the scalability limits and information-sharing constraints inherent to centralized design-time approaches [176]. These enforcers ensure correctness, bounded deviation, and completeness, which makes them well-suited for applications like collision avoidance and cooperative planning. Yet, static monitoring strategies remain insufficient against adaptive adversaries, which can evolve policies to bypass fixed defenses. To this end, adaptive monitoring frameworks such as Adversarial Markov Games cast detection as a dynamic interaction between attackers and defenders, showing that reinforcement learning can optimize both evasive attacks and responsive defenses in black-box settings [217]. These results demonstrate that in addition to runtime enforcement and decentralization, successful monitoring also necessitates the flexibility to adjust to adversarial co-evolution.
4.5 Standards and Organizational Measures
In addition to technical defenses, organizational frameworks and standards play a crucial role in shaping secure deployment of agentic AI in practice. These measures provide common guidelines, risk management practices, and reference architectures that organizations can adopt to prevent systemic vulnerabilities. For instance, the NIST AI Risk Management Framework (AI RMF) Generative AI Profile [5], developed under Executive Order 14110, provides a cross-sectoral reference for managing risks specific to generative and agentic AI. It extends the AI RMF by defining risks that are novel to or exacerbated by generative AI, such as the misuse of LLMs, unverified tool access, and autonomy-driven escalation. It outlines governance, mapping, measuring, and management practices to mitigate these risks. Complementary initiatives include the NCCoE Cyber AI Profile [157], which offers implementation guidance, Cybersecurity Framework, for integrating cybersecurity controls into AI systems; the OWASP Agentic AI Threats project [162], which catalogs common vulnerabilities such as insecure orchestration and prompt injection; and the CSA MAESTRO framework [54], which introduces a multi-layered threat modeling methodology tailored for agentic AI, providing structured analysis of risks across the AI lifecycle, from foundation models to multi-agent ecosystems. Together, these standards and organizational frameworks establish structural safeguards, embedding AI agents within broader ecosystems of regulation, compliance, and risk management.
Table 1: Secure-by-Design Defense Mechanisms Coverage Across Agentic AI Security Approaches.
Defense Approach Input Governance Tool-Call Guards Policy Engine Sandboxing/ Isolation Rate Limiting Audit Trails Rollback/ Recovery
Prompt-Injection-Resistant Designs
PromptArmor [197]
StruQ [34]
SecAlign [36]
Sign-Prompt [206]
Attn. Tracker [97]
JATMO [171]
TaskShield [103]
Melon [268]
Policy Filtering & Enforcement
LlamaGuard [99]
LlavaGuard [88]
GuardAgent [239]
AgentSpec [226]
ShieldAgent [41]
R²-Guard [110]
Sandboxing & Capability Confinement
SandboxEval [174]
IsolateGPT [238]
ToolEmu [188]
Monitoring & Adaptive Systems
Pro2Guard [225]
Arms Race [217]
Dec. RE [176]
CB [273]
Input Governance: Input validation, sanitization, and prompt injection detection mechanisms to prevent malicious inputs from reaching the agent. Tool-Call Guards: Runtime verification and constraints on tool/API invocations to prevent unauthorized or dangerous actions. Policy Engine: Formal policy specification and enforcement mechanisms to ensure agent behavior aligns with safety requirements. Sandboxing/Isolation: Isolated execution environments that contain potential damage from risky operations. Rate Limiting: Request throttling and resource usage controls to prevent denial-of-service and resource exhaustion. Audit Trails: Logging and traceability of agent actions for accountability and post-incident analysis. Rollback/Recovery: Mechanisms to reverse harmful actions or restore the system to safe states after detection of violations.
Annotations = Explicitly supported; = Not addressed or not applicable.
5 Benchmarks and Evaluation
To assess the impact of various agentic AI security vulnerabilities associated we discussed previously (as well the potency of defense strategies) it is necessary to undertake system evaluation via robust benchmarks. The first benchmarks proposed for agentic AI primarily focused on competence and sought to study whether an agent could complete a specified task under controlled conditions. Lately, as deployments edge closer to production, evaluation has shifted toward reliability, safety, and control. Consequently, recently DoomArena [19] has introduced security-oriented evaluation frameworks, a modular and configurable platform that enables cross-environment attack composition and fine-grained threat modeling for systematically probing agent failure modes and defense limitations. Thus, in this section, we cover: (i) landscape benchmarks used to gauge general capability, (ii) security–specific frameworks that probe failure modes such as harmful misuse, and (iii) methodological advances that improve fidelity, comparability, and reproducibility. Table 2 summarizes existing benchmarks spanning both capability and security-specific focuses, highlighting their domains, safety focus, and evaluation methodologies.
Table 2: Agentic AI Security and Capability Evaluation Benchmarks.
Benchmark Domain Safety Focus Process-Aware Multi-Turn Sandboxed Judge Type Reliability Metric Keypoints
Capability Benchmarks
BrowserGym [45] Web Rule-based End-state Unified interface
MiniWoB++ [133] Web Rule-based End-state Baseline
$\tau$-bench [248] Multi-domain Rule-based $\landk, pass$@k Consistency
GTA [227] Multi-domain Rule-based End-state, step-wise Tool agent
OSWorld [240] Computer Rule-based End-state, scripted Computer environment
OSWorld-Human [2] Computer Rule-based Weighted efficiency score Temporal performance
VisualWebArena [117] Web $\sim$ End-state, fuzzy Visually grounded
AndroidWorld [179] Android Rule-based End-state Dynamic parameterization
WebShop [247] E-commerce Rule-based End-state, Programmatic Reward Scalable
AgentBench [134] Multi-domain Rule-based End-state, step-wise Eight environments
WCA [150] Web $\sim$ End-state, fuzzy Labor-intensive "chores"
CRAB [242] Multi-domain Rule-based Graph-based Graph-based evaluation
AssistantBench [252] Open web $\sim$ Realistic & time-consuming
Security-Specific Benchmarks
ST-WebAgent-Bench [125] Enterprise CuP, Risk Policy compliance
AgentHarm [9] Open-domain Compliance Jailbreak + competence
OS-Harm [8] Computer LLM Safety, Accuracy Desktop effects
R-Judge [253] Meta-eval $\sim$ LLM Risk recognition Risk assessment
ToolEmu [188] Tool-use Side-effects Emulation
AgentDojo [61] Tool-use Rule-based Utillity, ASR Dynamic Framework
InjecAgent [258] Tool-use Rule-based ASR Indirect Prompt Injection
ASB [260] Multi-domain Rule-based ASR, RR, BP, PNA, NRP Multi-scenario
PrivacyLens [195] Personalized Comm $\sim$ LR, Helpfulness Privacy norm awareness
Agent-SafetyBench [264] Tool-use LLM Safety Score, Helpfulness Comprehensive
SAB [251] Embodied AI $\sim$ Rejection Rate, Risk Rate Embodied agents in simulated environment
CAIBench [191] Cybersecurity Rule-based $\sim$ Simultaneous offense and defense evaluation
WASP [69] Web $\sim$ ASR (intermediate & end-to-end), Utility Security by incompetence
SafeArena [219] Web $\sim$ TSR, RR, NSS, ARIA Deliberate misuse focused
OAS [222] Multi-domain $\sim$ FR, DR Multi-user dynamics
MCP-SafetyB [272] MCP Rule-based TSR, ASR Real MCP servers
MCP-SecB [246] MCP Rule-based ASR, RR, PSR Real MCP servers
MAGPIE [108] Multi-domain $\sim$ LR, TSR (Constraints), MR, Consensus Multi-agent
[…]
Full forms: Leakage Rate (LR), Misclassification Rate (MR), Refusal Rate (RR), Protection Success Rate (PSR), Task Success Rate (TSR), Failure Rate(FR), Disagreement Rate (DR), Normalized Safety Score (NSS), ARIA risk levels, Performance under No Attack (PNA), Benign performance (BP), and Net Resilient Performance (NRP).
5.1 Landscape benchmarks for capability
A number of realistic and interactive testbeds for web and computer-use agents have been developed. BrowserGym brings many such tasks under one consistent interface and scoring protocol (e.g., MiniWoB++, WorkArena, WebArena, among others), reducing fragmentation and enabling like-for-like comparisons across LLMs and agent designs [45]. In particular, BrowserGym emphasizes unified observation and action spaces and experiments showcase cross-model evidence that agent performance is sensitive to both model choices and the environments [45]. On the long-standing MiniWoB/MiniWoB++ benchmarks, proposed methods have continued to make progress [133, 76, 196].
[…]
5.2 Security-Specific and safety-specific benchmarks
Several benchmarks now also aim to evaluate agentic safety, i.e., risks arising from autonomous action, tool use, and long-horizon interaction, rather than from static chat completion. We discuss these next.
Web agent safety in enterprise contexts. ST-WebAgentBench is an online, enterprise-focused benchmark for testing whether web agents avoid unsafe actions (e.g., destructive operations in business systems) while pursuing goals [125]. Unlike legacy suites that only score end-task success, ST-WebAgentBench emphasizes trustworthiness under realistic web front ends (such as DevOps workflows, e-commerce, and enterprise CRM). The authors also propose novel evaluation metrics such as (1) Completion Under Policy (CuP) (task completions that adhere to acceptable policies) and (2) Risk Ratio (quantifies security breaches). In this manner, this benchmark highlights two distinct error classes that are useful for deployment: (i) task success/failure and (ii) unsafe/non-compliant decision-making, with the latter often being under-measured in prior work [125].
Open-domain harmful behavior and misuse. AgentHarm measures whether tool-using agents comply with harmful requests across multiple harm categories and whether jailbreaks preserve agentic competence (i.e., the ability to carry out multi-step harmful tool sequences once refusal is bypassed) [9]. AgentHarm relies on synthetic tools to safely simulate realistic actions (e.g., email, search, etc.) while evaluating safety during evaluation. Reported results indicate that simple jailbreak templates can markedly increase compliance while leaving task-execution ability largely intact, an especially concerning finding for the security of autonomous agents [9].
General computer-use safety. OS-Harm builds on OSWorld’s [240] full desktop environment to evaluate agent safety across office applications and file operations, scoring both accuracy and adherence to safety guidelines via LLM-based judges with validated agreement levels to human annotations [8]. Compared to pure web benchmarks, OS-Harm stresses desktop-level side-effects (e.g., unintentional data exfiltration or copyright infringement edits) and probes how trace format (screenshots and accessibility trees) affects automatic judging reliability [8].
Risk awareness and trace-level assessment. Rather than elicit harmful behavior, R-Judge assesses whether LLMs (and by extension, agents) can recognize safety risks in agent trajectories, providing a complementary lens for building better judge or guard(rail) components [253]. Meanwhile, ToolEmu evaluates agents within an LLM-emulated tool sandbox, enabling scalable probing of risky behaviors and potential negative side-effects [188]. The author’s findings via the emulator reveal that several LLM agents are prone to the aforementioned issues, thereby hindering real-world deployment.
Red-team corpora, dynamic web testbeds, and OS-agent sandboxes. Beyond the benchmarks discussed above, additional frameworks strengthen these evaluation dimensions. Red-team corpora like HarmBench [145] and JailbreakBench [29], though primarily LLM-focused, provide adversarial baselines applicable to agent-underlying models. Extending evaluation to realistic web interaction scenarios, dynamic web testbeds such as AgentDojo [61] and WASP [69] assess prompt injection vulnerabilities in web agents. OS-agent sandboxes including ToolEmu [188] and OS-Harm [8] further enable safe evaluation across desktop platforms.
5.3 Evolution of Agentic Security Evaluation
We now discuss some potential directions for the evolution of benchmarks being proposed for agentic AI security evaluation. Alongside insights based on current advances and progress, we discuss how new benchmarks can further augment evaluations by incorporating additional information or adopting relevant strategies.
Process-aware evaluations. End-state metrics (success/fail) miss important aspects of agent safety such as near-misses, unsafe tool invocations later rolled back, or brittle plans that succeed only stochastically. Newer benchmarks therefore score trajectory segments (plans, tool calls, and intermediate states) and use trace-level judges to detect policy violations or side-effects [260, 69, 125, 253, 188, 8]. The shift to process-aware evaluation unlocks finer-grained feedback for training guardrails and control measures, especially relevant for security-critical domains.
[…]
Standardizing judges and reducing judge bias. LLM-as-a-judge [83] is attractive for scale but can be biased by prompt design, trace format, or model choice. Several papers investigate how to structure judges (rubrics, multi-criteria prompts), validate them against humans, and reduce hallucinated assessments [8, 253]. An organic evolution of the judge approach, Agent-as-a-Judge, embeds an evaluative agent that reasons over trajectories and provides structured critique and scoring [270]. For safety-critical scenarios, when judges are used for evaluation, it is important to report: (i) inter-rater agreement with humans, (ii) sensitivity to trace redaction/format, as well as (iii) stability across random seeds and slight task paraphrases.
Sandboxing and emulation to contain risk. Security evaluation often requires testing unsafe prompts or tool sequences; executing these against live systems is risky and irreproducible [188, 8, 124, 219, 140]. Emulation (ToolEmu) and VM-backed sandboxes (OSWorld/OS-Harm) provide safe, deterministic environments for repeatable experiments [188, 8]. A desirable property is fidelity: the closer the emulator’s API/latency/error modes are to the real use-case, the more useful the results. In general, for security evaluations of agentic AI frameworks, it is especially important that upcoming benchmarks aim to report these fidelity assumptions explicitly.
Reproducibility and comparability. Recent meta-benchmarks (e.g. BrowserGym [45]) emphasize unified logging, seeded randomness, and fixed observation/action spaces to avoid apples-to-oranges comparisons. For security applications, future benchmarks should continue to provide: (i) public release of attack templates and defense configurations; (ii) reporting cost and latency (in particular safety systems that often add overhead); (iii) documenting environment determinism (web UIs and APIs can potentially be dynamic); and (iv) dataset hygiene to prevent contamination.
6 Open Challenges
We now discuss some open challenges in the field of agentic AI security. More specifically, we cover issues associated with measuring long-horizon safety in agentic systems, approaches for multi-agent system security, and developing better benchmarks for security evaluation of AI agents, among other considerations.
6.1 Long-Horizon Security
In general, LLM performance across long-horizon tasks (i.e. those that require planning, tool-use, sequential context management, and memory interactions across long episodes) suffers in comparison to shorter-term tasks [37, 129]. Recent computer use testbeds such as OSWorld [240] and OSWorld-Human [2] expose substantial performance and reliability gaps across long workflows and heterogeneous apps, even for strong agents. Similarly, agents struggle to manage memory appropriately over long horizons, leading to erroneous or redundant context [90]. This deficiency is only compounded for security considerations [205] as security risks can surface on multi-step tasks with tool use and partial observability that otherwise might not appear for shorter tasks. The open challenge thus is to measure and improve agentic AI safety across very long time/episode horizons. More specifically, progress can be made in two directions: (i) temporal robustness: whether agents can maintain safe behavior across multi-step subgoals, interruptions, and/or distribution shifts; and (ii) latent misbehavior detection: recent work [96] has shown that deceptive policies can persist through standard safety training and only activate under triggers, undermining short-episode audits.
6.2 Novel Multi-Agent Security Considerations
Multi-agent designs promise fault tolerance and specialization, but collaborations across agents also amplify attack surfaces and introduce novel security threats. Past work [92] has shown that even a single faulty or adversarial agent can cascade failures across different multi-agent organization topologies, and resilience varies sharply with organizational structure/hierarchy. Moreover, the presence of reviewers or challenge mechanisms [92] can improve error recovery significantly. Other work [87] has highlighted vulnerabilities in inter-agent communication protocols (such as message interception and manipulation attacks) and shown that these threats can compromise real-world agentic AI applications. Thus, future work can drive progress along the following directions: (1) robust messaging channels with authentication/verification that do not degrade performance (e.g. by introducing additional latency); (2) safety mechanisms for disagreement, contestation, and rollback that limit the potential for attacks caused by adversarial agents; and (3) developing sentinel agents that can better regulate the actions of worker agents, to safeguard the system against those that are malicious or faulty.
6.3 Improved Safety and Security Benchmarks
A more realistic evaluation of current safety vulnerabilities and deficiencies of agentic systems can only be undertaken if safety/security benchmarks can capture all possible attack scenarios and evaluate threat paradigms rigorously. There are many potential directions for future work to augment safety evaluations for agentic AI systems. (1) New distributional coverage metrics that score trajectory segments, as opposed to only considering the end-state (i.e. whether a task was completed successfully) and capture the entire distribution of performance (instead of average performance or a single run). (2) LLMs-as-a-judge are increasingly being used for agentic AI evaluation [270], but it is not clear how reliable or resilient the judge framework is, given that past work has shown LLM judges can suffer from critical security issues themselves [215]. Future work can thus work towards developing judges that are standardizes and robust to adversarial influence. (3) To truly assess how damaging potential attacks against agent systems can be, it is important to undertake rigorous testing in sandbox and emulated environments – future research work can aim to uncover whether or not current environments possess high fidelity with the real-world, and propose alternatives where current systems are inadequate.
6.4 Safety Against Adaptive Attacks
Adaptive attacks constitute a stronger threat model, where the attackers already have knowledge of the defense method employed by the system [257]. Recent studies show that adaptive attacks have been largely unexplored and ignored, as most work undertakes static (i.e., non-adaptive) defense evaluations for AI agents [257, 105, 97]. More generally, in adversarial attack research in ML/AI, many defenses initially reported as effective are commonly undermined soon after their release, primarily because they lack thorough evaluation against adaptive attack strategies [105, 11, 22]. In the same vein, although different studies are being done to consider, defend, and evaluate adaptive attacks [197, 217], their potential impacts across agentic AI systems merit further exploration. Hence, future works can can explore the following directions: (i) proposing and defining novel evaluation methods specific to adaptive attacks, and (ii) exploring the impact of current defenses against adaptive attacks, and (iii) developing stronger adaptive defenses that can co-evolve and thwart adaptive attackers.
6.5 Agentic AI Security in the Physical World
While this survey primarily focuses on software-based agents, physically embodied agents introduce additional and largely unexplored security risks due to their dependence on hardware and direct interaction with the physical environment. (1) Such risks involve adversaries manipulating robot perception through sensor spoofing, including GPS spoofing that redirects vehicles into restricted zones [114, 213], LiDAR laser projection spoofing that injects phantom obstacles triggering emergency braking [21, 200], and adversarial patches on road signs and persons causing threats through misclassification [70, 212]. (2) Ensuring robust and secure operation of physically embodied agents remains a critical challenge, as agent architecture increasingly determines vulnerability. End-to-end foundation models (e.g., RT-2 [271], OpenVLA [115]) map raw sensory inputs directly to motor actions, eliminating defensive separation and prioritizing zero-shot generalization, but without corresponding improvements in adversarial robustness [250]. Addressing this trade-off between generalization and safety, requires new methods to preserve internal defenses and safeguard agents against real-world sensor and hardware attacks. (3) Beyond mobile robotics, LLM-based agents are increasingly deployed in smart home environments for device control and automation [186], yet systematic security evaluation and benchmark in these environments remains underexplored.
6.6 Human-Agent Security Interfaces
The boundary between humans and agentic AI systems introduces a unique security frontier. On one hand, agents are often deployed to interact directly with end-users, who may provide instructions, corrections, or verification of system behavior. On the other hand, human oversight itself can be adversarially influenced. For instance, users can be socially engineered by an attacker into approving unsafe actions [^1] Consider an example from [81] where the attacker lures the user by saying: “you won’t believe ChatGPT’s response to this prompt!” followed by the prompt injection text in another language or Base64 so the user cannot ascertain its adversarial nature., or become overwhelmed by the complexity of long execution traces. These potential vulnerabilities highlight the need to better understand and secure the interface layer between agents and human operators. Thus, future work can investigate several problems to bridge this gap, such as: (i) designing user interaction mechanisms that are robust to manipulation and do not overly burden the human with verification tasks; (ii) developing auditing and summarization tools that make long-horizon agent behavior interpretable enough for reliable user verification; and (iii) studying how attackers can exploit user trust, particularly in multi-step workflows where human approval serves as a safety check. We posit that progress on this front will require not only technical defenses but also empirical studies on human factors, as the security of agentic systems depends as both human susceptibility as well as algorithmic robustness.
7 Conclusion
In this survey, we provided an in-depth analysis of current work in agentic AI security. Our paper first discussed the unique landscape of security threats that agentic AI systems are susceptible to via a comprehensive taxonomy of related work (Section 3). Then, we have discussed several defense strategies and security controls that can be employed to mitigate known attack vectors (Section 4) as well as various benchmarks and evaluation metrics to guide rigorous testing of proposed agentic attack/defense approaches (Section 5). Finally, we discussed a number of open challenges where progress can be made, and how doing so will significantly augment the safety properties of future agentic AI systems (Section 6). Our goal through this survey article is to add to the existing body of work on agentic AI security by providing a distilled introduction to the field, and to galvanize research progress in making agentic frameworks more safe and secure for large-scale societal use.